# OpenGait: A Comprehensive Benchmark Study for Gait Recognition towards Better Practicality

### v1

@@ VII Conclusion and Future Work
This work embarks on a journey to fulfill the urgent need for practical gait recognition by basic codebase building, previous SoTA revisiting, and new baseline construction, exhibiting a comprehensive benchmark study aimed at catalyzing further advancements in this field. As a result, OpenGait serves as an easy-to-use platform, while DeepGaitV2, SkeletonGait, and SkeletonGait++ are proposed as representatives of the appearance-based, model-based, and multi-modal gait recognition baseline models, respectively.
Beyond the above contributions, this work also makes it clear that the dataset scale, model size, and modality designs stand out as three of the bottlenecks in gait recognition. While these aspects are commonly acknowledged in computer vision, our experiments offer concrete insights into why the gait community continues to rely predominantly on shallow gait models and binary silhouette images. This work presents a good attempt to expose these gait-specific issues, recognizing that there remains considerable ground to cover for real-world gait recognition applications.
VIII Acknowledgement
We would like to thank …

### v2

@@ VIII Conclusions
With gait recognition becoming increasingly practicable, this work addresses the urgent need for developing a foundational gait codebase, revisiting previous state-of-the-art methods, and establishing new baselines. We provide a comprehensive benchmark study to encourage further advancements in the field. OpenGait offers an accessible platform, while DeepGaitV2, SkeletonGait, and SkeletonGait++ represent new baseline models for appearance-based, model-based, and multi-modal gait recognition, respectively.
Moreover, our work highlights three main challenges in gait recognition: dataset scale, model size, and modality design. Typically, indoor settings facilitate obtaining permission from volunteers for gait research but struggle to replicate real-world conditions. Conversely, outdoor gait datasets better capture real-world changes but present challenges in acquiring the walking crowd with consent. Therefore, the research community must reduce its heavy reliance on large-scale datasets. Emerging unsupervised and self-supervised methods offer promising solutions. Building upon the dataset issues, our comprehensive experimental investigation addresses the problem of determining suitable capacity for deep gait models, exemplified by the DeepGaitV2 series. Beyond only the appearance-based gait methodology, this paper further reveals the continued need for a discriminative yet clean gait modality by systematically comparing three proposed baseline models. By highlighting these challenges and discussing emerging trends and opportunities, we hope to pave the way for future research aimed at improving real-world gait recognition applications.
IX Acknowledgements
We sincerely express our heartfelt gratitude to the OpenGait community for their invaluable support and collaboration. The insights provided by the Associate Editor and Reviewers of this paper have been instrumental in shaping Sec. VII, for which we are deeply thankful. We extend our special thanks to Mr. Jilong Wang, Mr. Beibei Lin, Mrs. Yunjie Peng, Mr. Zirui Zhou, and Mr. Dingqiang Ye for their significant contributions of code to OpenGait. We also deeply appreciate the support and extensions provided by the teams behind Gait3D [9], GREW [8], FastPoseGait [12], CASIA-E [81], CCPG [20], and CCGR [23], which have greatly enriched the field and our work.
This work is jointly supported by the National Natural Science Foundation of China (62476120, 62206022, 62276025, 62476027), Shenzhen International Research Cooperation Project (GJHZ20220913142611021), Beijing Municipal Science & Technology Commission (Z231100007423015), and Scientific Foundation for Youth Scholars of Shenzhen University (868-000001033383).
We would like to thank Mr. Chunhua Wang for his professional technical support to our work.